# 03F40 (primary), 03D20

### v2

@@ 6 Fundamental Sequences
[…]
Fundamental sequences are sequences $(\xi[n])_{n\in\mathbb{N}}$, such that when $\xi$ is a limit, $\xi[n]\to\xi$. They are defined for ordinals written in $\psi$-normal form.
@@
The reader may assume that all expressions above are in $\Omega$}-normal form as per Definition 4.1, but actually the value of $\tau(\xi)$ will not be sensitive to how $\xi$ is written. Note that successors are assigned cofinality $\omega$, so this is not a ‘true’ cofinality, but this definition is convenient for the fundamental sequences below.
@@
$\psi[\varepsilon_{{\Omega}+1}]=\{\xi<\varepsilon_{{\Omega}+1}:\xi^{*}<1+\psi(\varepsilon_{{\Omega}+1})\}.$
@@
  8. 8.
    if $\tau(\zeta)={\Omega}$ and $\xi=\psi(\zeta)$, then $\xi[0]=0$
    and $\xi[\theta+1]=\psi(\zeta[\xi[\theta]])$,
where all expressions are written in either $\Omega$}-decomposition or $\psi$-normal form.
These fundamental sequences converge and preserve $\psi$-normal forms (see e.g. [19]).
@@
Moreover, if $\psi(\zeta)$ is in $\psi$-normal form and $i>0$, then:
  1. 1.
@@
We also note that fundamental sequences mostly map uncountable ordinals to uncountable ones.
@@
Let $\zeta\in\psi[\varepsilon_{{\Omega}+1}]$ and $\iota<\tau(\zeta)$. If $\zeta>{\Omega}$, then either $\zeta[\iota]\in\{0,1\}$ and $\iota=0$, or else $\zeta[\iota]$ is uncountable.
Proof 6.7.
Write $\zeta={\Omega}^{\eta}{\alpha}+\gamma$ and proceed by induction on $\zeta$. Since $\zeta>{\Omega}$, $\zeta[\iota]$ is countable only if $\gamma=0$. If $\alpha$} is a limit, then $\zeta[\iota]={\Omega}^{\eta}({\alpha}[\iota])$, which is either uncountable or zero. If $\alpha$} is a successor, then $\zeta[\iota]$ is uncountable.
Otherwise, ${\alpha}=1$. If $\eta$ is a limit, then $\zeta[\iota]$ can only be countable when $\eta[\iota]=0$, whence $\zeta[\iota]={\Omega}^{0}=1$. Otherwise, $\eta=\delta+1$ is a successor and $\delta>0$ since $\zeta>{\Omega}$, so $\zeta[\iota]$ is uncountable or zero when $\iota=0$.
@@
Here, the expression $y=f(x)$ should be represented according to the conventions of Theorem 2.17.
@@
Standard proof-theoretic analyses are based on variants of $\psi$ which take only additively indecomposable values (or even strongly critical ones), so there is some verification involved in applying Theorem 6.9 to our ‘slow’ $\psi$. Let $\psi^{+}$ denote such a collapsing function, so that $\psi^{+}(0)=1$ and, for $\xi>0$,
$\psi^{+}(\xi)=\sup\{\omega^{\psi^{+}(\zeta)+1}:\zeta<\xi\text{ and }\zeta^{*}<\psi^{+}(\zeta)\}.$
In view of results of Cichon et al. [7], we only require that $\psi(\varepsilon_{{\Omega}+1})=\psi^{+}(\varepsilon_{{\Omega}+1})$ and that our fundamental sequences satisfy certain regularity properties. Verification of these facts can be performed by noting that the desired equality follows from the identity $\psi^{+}(\alpha)=\psi(\Omega\alpha)$ and the proofs of regularity in e.g. [19], specialize to our more restrictive $\psi$. The latter applies also to Lemma 6.5.
7 Independence
Given Theorem 6.9, in order to to prove independence of Theorem 2.16, it suffices to compare our Goodstein process to the process of stepping down the fundamental sequences. As pointed out above, the ordinal interpretation $o_{B}$ is not suitable here. Instead, we will introduce a second ordinal interpretation, $u_{B}$, based on $\psi$. The issue here is that $\psi$ does not have some of the key properties of $\vartheta$ which allowed us to define a monotone ordinal assignment, and thus $u_{B}$ will not be monotone. This, however, will not be a problem.
To illustrate this, let us turn to the proof of independence of the original Goodstein principle. In this setting, one typically shows that for $b\geq 2$,
@@ 7 Independence
Suppose that we show that for every $n$ and $b\geq 2$, there is $n^{\prime}<\left(\begin{smallmatrix}b+1\\ b\end{smallmatrix}\right)n$ such that $\left(\begin{smallmatrix}\omega\\ b+1\end{smallmatrix}\right)n^{\prime}=\left(\begin{smallmatrix}\omega\\ b\end{smallmatrix}\right)n[b-1]$. For example, if $b=3$ and $n=3^{3}$, we could take $n^{\prime}=4^{3}$, and observe that
$\left(\begin{smallmatrix}\omega\\ 4\end{smallmatrix}\right)n^{\prime}=\omega^{3}=\omega^{\omega}[2]=\left(\begin{smallmatrix}\omega\\ 3\end{smallmatrix}\right)n[2].$
@@
Thus our function $u_{B}$ will be such that, letting $u_{i}=u_{\mathcal{O}_{i}}$ and ${\uparrow}_{i}={\uparrow}_{\mathcal{O}_{i}}^{\mathcal{O}_{i+1}}$, for every $n\in\mathbb{N}$ there is $n^{\prime}<{\uparrow}_{i}n$ such that $u_{i+1}(n^{\prime})=u_{i}(n)[i+1]$. We can then replicate the above argument to show that the Goodstein process majorizes the function $F$ of Definition 6.8. The function $u_{B}$ follows a similar pattern to $o_{B}$, where a ‘large’ ordinal assignment $U_{B}$ is first defined, and then ‘collapsed’ to $u_{B}$.
@@
Despite being simpler to work with in many ways, the main drawback of the $\psi$ function over $\vartheta$ is the need to check whether ordinals are in normal form before applying standard operations to them. Likewise, the $u_{B}$ and $U_{B}$ functions come with normal forms that allow them to work as intended, despite the lack of monotonicity.
@@
Proof 7.4.
Let $b={\rm Base}_{B}(n)$. The $U_{B}$-normal form condition is unchanged from the original definition and is not hard to check that $(U_{B}(n))^{*}=\max u_{B}[{\rm Digit}_{b}(n)]$. Thus, $(U_{B}(n))^{*}<\psi(U_{B}(n))$ if and only if for every $c\in{\rm Digit}_{b}(n)$, $u_{B}(c)<u_{B}(n)$.
@@
Next we see that $U_{B}(6)=U_{B}^{6}(6)={\Omega}$, so $u_{B}(6)=\psi({\Omega})=u_{B}(2)$. This already shows that $u_{B}$ is not an injective function. In fact it is not even non-decreasing, as $u_{B}(4)>u_{B}(6)$.
@@
If moreover $n$ is in $U^{b}_{B}$-normal form and $y\in{\rm Digit}_{c}(\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n)$, then by Lemma 3.83, $y={\uparrow}x$ for some $x\in{\rm Digit}_{b}(n)$, which by assumption implies that $y$ is in $u_{C}$-normal form. Moreover, $e$ and $r$ are in $U^{b}_{B}$-normal form, which by the induction hypothesis implies that $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e$ and $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle r$ are in $U^{c}_{C}$-normal form. Finally, ${\Omega}^{U^{c}_{C}(\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e)}u_{C}({\uparrow}a)+U^{c}_{C}(\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle r)$ is in $\Omega$}-decomposition since ${\Omega}^{U^{b}_{C}(e)}u_{B}(a)+U^{b}_{B}(r)$ is in $\Omega$}-decomposition and the respective ordinals are equal.
@@
With transformations from $B$ to a good successor $C$, for all $n\in\mathbb{N}$, $u_{B}(n)=u_{C}({\uparrow}n)$ and $U_{B}(n)=U_{C}({\uparrow}n)$. Moreover, if $n$ is in $u_{B}$-normal form, then ${\uparrow}n$ is in $u_{C}$-normal form, and if $n$ is in $U_{B}$-normal form, then ${\uparrow}n$ is in $C_{C}$-normal form,
Proof 7.9.
Proceed to prove all claims by induction on $n$. We first prove that $U_{B}(n)=U_{C}({\uparrow}n)$. Let $b={\rm Base}_{B}(n)$ and $c={\rm Base}_{C}({\uparrow}n)$. By Lemma 3.42, ${\uparrow}n=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n$. The induction hypothesis yields that for $x\in{\rm Digit}_{b}(n)$, $u_{B}(x)=u_{C}({\uparrow}x)$, and if $x$ is in $u_{B}$-normal form, then ${\uparrow}x$ is in $u_{C}$-normal form. Lemma 7.6 then yields $U_{B}(n)=U^{b}_{B}(n)=U^{c}_{C}({\uparrow}n)=U_{C}({\uparrow}n)$. From the condition $U_{B}(n)=U_{C}({\uparrow}n)$, we then obtain
@@
Now, assume that $n$ is in $u_{B}$-normal form. Then, $n$ is in $U^{b}_{B}$-normal form, so by Lemma 7.6, ${\uparrow}n=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n$ is in $U^{c}_{C}$-normal form, which since $c={\rm Base}_{C}({\uparrow}n)$ implies that ${\uparrow}n$ is in $U_{C}$-normal form. Let $y\in{\rm Digit}_{c}({\uparrow}n)$. By Lemma 3.83, $y={\uparrow}x$ for some $x\in{\rm Digit}_{b}(n)$. Since $n$ is in $u_{B}$-normal form, $x$ is in $u_{B}$-normal form and $u_{B}(x)<u_{B}(n)$. By the induction hypothesis, $y={\uparrow}x$ is in $u_{C}$-normal form and
@@
where the first equality is by the induction hypothesis and we have already established the second.
Next, our goal is to match up the fundamental sequences with values of the new ordinal interpretation. We begin with the uncountable part.
@@
Let $\iota<\tau(U_{B}(s))$ be such that $\iota=u_{B}(c)$ for some $c<b$ in $u_{B}$-normal form, and either $\iota$ is infinite, or for every $b$-digit $a$ of $s$, there is $a^{\prime}<a$ in $u_{B}$-normal form such that $u_{B}(a^{\prime})=u_{B}(a)[\iota]$.
Then, there is $s^{\prime}<s$ in $U^{b}_{B}$-normal form such that $U(s^{\prime})=U(s)[\iota]$.
Proof 7.11.
Proceed by induction on $s$. If $s<b$, then $U(s)=u_{B}(s)<{\Omega}$ and thus has countable cofinality, so $\iota$ is finite. By assumption, we can find $s^{\prime}<s$ in $u_{B}$-normal form (which, since $s^{\prime}<b$, implies that $s^{\prime}$ is in $U^{b}_{B}$-normal form by definition) with $u_{B}(s^{\prime})=u_{B}(s)[\iota]$, and since $s^{\prime}<s<b$, we have that $U(s^{\prime})=u_{B}(s^{\prime})$, so

### v4

@@ 6. Fundamental Sequences
From the definition, one obtains $\psi(n)=n+1$ for $n<\omega$ and $\psi(\omega)=\omega$. Since $\omega^{*}=\omega=\psi(\omega)$, $\psi(\omega)$ does not affect the computation of $\psi(\omega+1)$, and thus we also have $\psi(\omega+1)=\omega$. Similarly, we have that $\psi({\alpha})=\omega$ whenever $\omega\leq{\alpha}<{\Omega}$, and thus $\psi({\Omega})=\omega$ as well. In general, $\psi$ is non-decreasing, but not strictly increasing. However, we obtain an increasing function if we restrict the domain to ordinals in $\psi$-normal form.
@@
[…]
In order to connect ordinal notation systems to $\Sigma^{0}_{1}$ independence, we may equip them with fundamental sequences, i.e. families of sequences $(\xi[n])_{n\in\mathbb{N}}$, such that when $\xi$ is a limit, $\xi[n]\to\xi$. Such sequences exist for countable limits, as well as some uncountable ordinals with countable cofinality.
@@
Note that successors are assigned cofinality $\omega$, so this is not a ‘true’ cofinality, but this definition is convenient to uniformly define fundamental sequences. It is convenient to also define fundamental sequences for ordinals $\xi$ of cofinality $\Omega$}, but in this case, $\xi[\theta]$ must be defined for all $\theta<{\Omega}$.
@@
$\psi[\varepsilon_{{\Omega}+1}]=\{\xi<\varepsilon_{{\Omega}+1}:\xi^{*}<\psi(\varepsilon_{{\Omega}+1})\}.$
@@
  8. (8)
    if $\tau(\zeta)={\Omega}$ and $\xi=\psi(\zeta)$, then $\xi[0]=0$ and $\xi[\theta+1]=\psi(\zeta[\xi[\theta]])$,
where all expressions are written in either $\Omega$}-decomposition or $\psi$-normal form, accordingly.
As a useful example, it is readily checked that $\omega[i]=i$ for all $i$, using the fact that $\omega=\psi({\Omega})$. These fundamental sequences converge and preserve $\psi$-normal forms (see e.g. [20]).
@@
Moreover, if $\psi(\zeta)$ is in $\psi$-normal form and $i<\omega$, then:
  1. (1)
    $\psi(\zeta)$ is a successor if and only if $\zeta$ is a successor, in which case $\psi(\zeta)=\psi(\zeta[i])+1$ and $\zeta=\zeta[i]+1$.
  2. (2)
@@
We also note that fundamental sequences ‘almost always’ map uncountable ordinals to uncountable ones.
@@
Let $\zeta\in\psi[\varepsilon_{{\Omega}+1}]$ and $\iota<\tau(\zeta)$.
  1. (1)
    If $\zeta>{\Omega}$, then either $\zeta[\iota]\in\{0,1\}$ and $\iota=0$, or else $\zeta[\iota]$ is uncountable.
  2. (2)
    If $\zeta$ is a limit and $\iota=\omega\lambda+\ell$ with $\ell<\omega$, then $\zeta[\iota]$ is of the form $\omega\rho+r$ with $r\leq\max\{2,\ell\}$.
Proof.
(1) Write $\zeta={\Omega}^{\eta}{\alpha}+\gamma$ and proceed by induction on $\zeta$. Since $\zeta>{\Omega}$, $\zeta[\iota]$ is countable only if $\gamma=0$. If $\alpha$} is a limit, then $\zeta[\iota]={\Omega}^{\eta}({\alpha}[\iota])$, which is either uncountable or possibly zero with $\iota=0$. If $\alpha$} is a successor, then $\zeta[\iota]$ is uncountable.
Otherwise, ${\alpha}=1$. If $\eta$ is a limit, then $\zeta[\iota]$ can only be countable when $\eta[\iota]=0$, whence $\iota=0$ and $\zeta[\iota]={\Omega}^{0}=1$. Otherwise, $\eta=\delta+1$ is a successor and $\delta>0$ since $\zeta>{\Omega}$, so $\zeta[\iota]$ is uncountable or zero when $\iota=0$.
(2) Proceed by induction on $(\iota,\zeta^{*},\zeta)$, ordered lexicographically (i.e., a primary induction on $\iota$, with a secondary induction on $\zeta^{*}$, and a tertiary induction on $\zeta$).
  1. Case 1
    ($\zeta<{\Omega}$). Write $\zeta=\psi(\xi)$ in $\psi$-normal form, so that $\xi^{*}<\zeta=\zeta^{*}$. In this case, $\tau(\zeta)=\omega$, so $\iota$ is finite and thus we must have $\iota=\ell$. Since $\psi(\xi)$ is a limit in $\psi$-normal form, $\xi$ is uncountable; for, if $\xi$ is finite, then $\psi(\xi)=\xi+1$, which is not a limit, and if $\xi$ is countably infinite, then $\psi(\xi)=\omega\leq\xi$, so $\psi(\xi)$ cannot be in $\psi$-normal form. Consider the following sub-cases.
    1. Case 1.1
      ($\xi={\Omega}$). We directly compute $\zeta[\iota]=\iota$, which is of the desired form since $\iota=\ell$.
    2. Case 1.2
      ($\xi>{\Omega}$). We have that $\zeta[\iota]=\psi(\xi[\iota^{\prime}])$, where $\iota^{\prime}$ depends on the cofinality of $\xi$. By (1), either $\xi[\iota^{\prime}]\leq 1$ or $\xi[\iota^{\prime}]$ is uncountable. In the first case, $\zeta[\iota]\leq\psi(1)=2$, so we may assume that $\xi[\iota^{\prime}]$ is uncountable. We consider two further sub-cases, according to the cofinality of $\xi$.
      1. Case 1.2.1
        ($\tau(\xi)=\omega$). In this case, $\iota^{\prime}=\iota$. Since $\xi^{*}<\zeta=\zeta^{*}$, we may apply the induction hypothesis to obtain $\xi[\iota]=\omega\rho+r$ with $r\leq\max\{2,\ell\}$ and $\rho>0$ (as we are assuming that $\xi[\iota]$ is uncountable). Then, using Lemma 6.5(1), we see that $\zeta[\iota]=\psi(\xi[\iota])=\psi(\omega\lambda+r)=\psi(\omega\lambda)+r$, which is of the desired form since $\psi(\omega\lambda)$ is a limit.
      2. Case 1.2.2
        ($\tau(\xi)={\Omega}$). Either $\iota=0$ and $\zeta[\iota]=0$, which is of the desired form, or $\zeta[\iota]=\psi(\xi[\iota^{\prime}])$, where $\iota^{\prime}=0$ if $\iota=0$ and otherwise $\iota^{\prime}=\zeta[\iota-1]$. If $\iota^{\prime}=0$, we may set $\rho^{\prime}=r^{\prime}=0$ and note that $\iota^{\prime}=\omega\rho^{\prime}+r^{\prime}$ with $r^{\prime}<\max\{2,\ell\}$. If $\iota^{\prime}=\zeta[\iota-1]$ the induction hypothesis is available since $\iota-1<\iota$ to obtain $\zeta[\iota-1]=\omega\rho^{\prime}+r^{\prime}$ with $r^{\prime}\leq\max\{2,\iota-1\}=\max\{2,\ell-1\}$. In either case, $\iota^{\prime}=\omega\rho^{\prime}+r^{\prime}$ with $r^{\prime}\leq\max\{2,\ell\}$. Applying the induction hypothesis once again, this time by $\xi^{*}<\zeta^{*}$, $\xi[\iota^{\prime}]=\xi[\omega\rho^{\prime}+r^{\prime}]=\omega\rho+r$ for some $r\leq\max\{2,r^{\prime}\}\leq\max\{2,\ell\}$. Since we are assuming that $\xi[\iota^{\prime}]$ is uncountable, $\rho>0$. Using Lemma 6.5(1), we see that $\zeta[\iota]=\psi(\xi[\iota^{\prime}])=\psi(\omega\rho+r)=\psi(\omega\rho)+r$, which is of the desired form.
  2. Case 2
    ($\zeta\geq{\Omega}$). Write $\zeta={\Omega}^{\eta}{\alpha}+\gamma$. If $\gamma>0$, the induction hypothesis applies readily to $\gamma$. Otherwise, $\zeta={\Omega}^{\eta}{\alpha}$. We cannot have $\zeta[\iota]$ be a successor if $\alpha$} is a limit, so we write ${\alpha}=\delta+1$ and see that $\zeta[\iota]={\Omega}^{\eta}\delta+{\Omega}^{\eta}[\iota]$, which is a successor if and only if ${\Omega}^{\eta}[\iota]$ is a successor. If $\eta=\rho+1$, then ${\Omega}^{\eta}[\iota]={\Omega}^{\rho}\iota$, which can only be a successor if $\rho=0$ and $\iota<\omega$, in which case we obtain $\zeta[\iota]={\Omega}^{\eta}\delta+\iota$, which is of the required form. If $\eta$ is a limit, ${\Omega}^{\eta}[\iota]={\Omega}^{\eta[\iota]}$, which can only be a successor if $\iota=0$ and $\eta[\iota]=0$, in which case $\zeta[\iota]={\Omega}^{\eta}\delta+1$, again of the desired form.∎
These fundamental sequences can be used to define very fast-growing functions, such as the function $F$ below.
@@ 2. Case 2
Here, the expression $y=f(x)$ should be represented according to the conventions of Theorem 2.
@@
Standard proof-theoretic analyses are based on variants of $\psi$ which take only additively indecomposable values (or even strongly critical ones), so there is some verification involved in applying Theorem 6.8 to our ‘slow’ $\psi$. Using the notation of Buchholz [5], let $\psi^{\mathbb{H}}$ denote the collapsing function which takes additively indecomposable values. According to Theorem 1 of Cichon et al. [7], we only require that $\psi(\varepsilon_{{\Omega}+1})=\psi^{\mathbb{H}}(\varepsilon_{{\Omega}+1})$ and that our fundamental sequences satisfy certain regularity properties. The desired equality can be verified using the identity $\psi^{\mathbb{H}}(\alpha)=\psi(\Omega\alpha)$ and the proofs of regularity in e.g. [20] specialize to our more restrictive $\psi$. Lemma 6.5 is an example of this specialization.
7. Independence
Given Theorem 6.8, in order to to prove independence of Theorem 1, it suffices to compare our Goodstein process to the process of stepping down the fundamental sequences. As pointed out above, the ordinal interpretation $o_{B}$ is not suitable here. Instead, we will introduce a second ordinal interpretation, $u_{B}$, based on $\psi$. The issue here is that $\psi$ does not have some of the key properties of $\vartheta$ which allowed us to define a monotone ordinal assignment, and thus $u_{B}$ will not be monotone. This makes $u_{B}$ unsuitable for proving termination of the fractal Goodstein process yet, surprisingly, it will not be a problem for proving independence.
To illustrate this, let us turn to the proof of independence of the original Goodstein principle. In this setting, one typically shows that, for $b\geq 2$,
@@ 7. Independence
Suppose that we show that for every $n$ and $b\geq 2$, there is $n^{\prime}<\left(\begin{smallmatrix}b+1\\ b\end{smallmatrix}\right)n$ such that $\left(\begin{smallmatrix}\omega\\ b+1\end{smallmatrix}\right)n^{\prime}=\left(\begin{smallmatrix}\omega\\ b\end{smallmatrix}\right)n[b-1]$. For example, if $b=3$ and $n=3^{3}$, we could take $n^{\prime}=4^{2}$, and observe that
$\left(\begin{smallmatrix}\omega\\ 4\end{smallmatrix}\right)n^{\prime}=\omega^{2}=\omega^{\omega}[2]=\left(\begin{smallmatrix}\omega\\ 3\end{smallmatrix}\right)n[2].$
@@
Thus our function $u_{B}$ will be such that, letting $u_{i}=u_{\mathcal{D}_{i}}$ and ${\uparrow}^{i+j}_{i}={\uparrow}_{\mathcal{D}_{i}}^{\mathcal{D}_{i+j}}$, for every $n\in\mathbb{N}$, there is $n^{\prime}<{\uparrow}^{i+j}_{i}n$ such that $u_{i+j}(n^{\prime})=u_{i}(n)[j]$. We can then replicate a variant of the above argument to show that the Goodstein process majorizes the function $F$ of Definition 6.7. The function $u_{B}$ follows a similar pattern to $o_{B}$, where a ‘large’ ordinal assignment $U_{B}$ is first defined, and then ‘collapsed’ to $u_{B}$.
@@
We can deduce some basic properties of $n$ from $U_{B}(n)$.
Lemma 7.2.
Let $B$ be a base hierarchy and $n\geq\min B$.
  1. (1)
[…]
  2. (2)
    If $\tau(U^{b}_{B}(n))={\Omega}$, then ${\rm Base}_{B}(n)\mid n$.
Proof.
(1) The equivalence of $u_{B}(n)$ and $U_{B}(n)$ being limits is a consequence of Lemma 6.5(1), so we check that $U_{B}(n)$ is a limit if and only if $\min B\mid n$. Let $b={\rm Base}_{B}(n)$ and $U=U^{b}_{B}$ and write $n=_{b}b^{e}a+r$, so that $U_{B}(n)={\Omega}^{U(e)}u_{B}(a)+U(r)$. Since $\min B\mid b$, we have that if $r=0$, then $\min B\mid n$ and $U_{B}(n)$ is a limit. If $0<r<\min B$, then $\min B\nmid n$ and $U(r)=r$, so $U_{B}(n)$ is not a limit. Otherwise, by induction, we have that $U(r)$ is a limit if and only if $\min B\mid r$, hence $U_{B}(r)$ is a limit if and only if $\min B\mid n$.
(2) If $\tau(U_{B}(n))={\Omega}$, then ${\Omega}\mid U_{B}(n)$, so by Lemma 4.5, ${\rm Base}_{B}(n)\mid n$.∎
Remark 7.3.
Note that (2) is not an equivalence, since for example if $b={\rm Base}_{B}(n)$ and $n=_{b}b^{1}a$ with $u_{B}(a)$ a limit, then $\tau(U_{B}(n))=\omega$, yet $b\mid n$. However, we will not need to precisely identify those $n$ with $\tau(U_{B}(n))={\Omega}$, and the condition ${\rm Base}_{B}(n)\mid n$ is a reasonable approximation.
Despite being simpler to work with in many ways, the main drawback of the $\psi$ function over $\vartheta$ is the need to check whether ordinals are in normal form before performing computations with them. Likewise, the $u_{B}$ and $U_{B}$ functions come with normal forms that allow them to work as intended, despite the lack of monotonicity.
@@
Proof.
Let $b={\rm Base}_{B}(n)$. The $U_{B}$-normal form condition is unchanged from the original definition and is not hard to check that $(U_{B}(n))^{*}=\max u_{B}[{\rm Digit}_{b}(n)]$. Thus, $(U_{B}(n))^{*}<\psi(U_{B}(n))$ if and only if for every $c\in{\rm Digit}_{b}(n)$, $u_{B}(c)<u_{B}(n)$. ∎
@@
Next, we see that $U_{B}(6)=U_{B}^{6}(6)={\Omega}$, so $u_{B}(6)=\psi({\Omega})=u_{B}(2)$. This already shows that $u_{B}$ is not an injective function. In fact it is not even non-decreasing, as $u_{B}(4)>u_{B}(6)$.
@@
If moreover $n$ is in $U^{b}_{B}$-normal form and $y\in{\rm Digit}_{c}(\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n)$, then by Lemma 3.5(3), $y={\uparrow}x$ for some $x\in{\rm Digit}_{b}(n)$, which by assumption implies that $y$ is in $u_{C}$-normal form. Moreover, $e$ and $r$ are in $U^{b}_{B}$-normal form, which by the induction hypothesis implies that $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e$ and $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle r$ are in $U^{c}_{C}$-normal form. Finally, ${\Omega}^{U^{c}_{C}(\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e)}u_{C}({\uparrow}a)+U^{c}_{C}(\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle r)$ is in $\Omega$}-decomposition since ${\Omega}^{U^{b}_{C}(e)}u_{B}(a)+U^{b}_{B}(r)$ is in $\Omega$}-decomposition and the respective ordinals are equal. ∎
@@
With transformations from $B$ to a good successor $C$, for all $n\in\mathbb{N}$, $u_{B}(n)=u_{C}({\uparrow}n)$ and $U_{B}(n)=U_{C}({\uparrow}n)$. Moreover, if $n$ is in $u_{B}$-normal form, then ${\uparrow}n$ is in $u_{C}$-normal form, and if $n$ is in $U_{B}$-normal form, then ${\uparrow}n$ is in $U_{C}$-normal form.
Proof.
Proceed to prove all claims by induction on $n$. We first prove that $U_{B}(n)=U_{C}({\uparrow}n)$. Let $b={\rm Base}_{B}(n)$ and $c={\rm Base}_{C}({\uparrow}n)$. By Lemma 3.3(2), ${\uparrow}n=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n$. The induction hypothesis yields that for $x\in{\rm Digit}_{b}(n)$, $u_{B}(x)=u_{C}({\uparrow}x)$, and if $x$ is in $u_{B}$-normal form, then ${\uparrow}x$ is in $u_{C}$-normal form. Lemma 7.7 then yields $U_{B}(n)=U^{b}_{B}(n)=U^{c}_{C}({\uparrow}n)=U_{C}({\uparrow}n)$. From the condition $U_{B}(n)=U_{C}({\uparrow}n)$, we then obtain
@@
Now, assume that $n$ is in $u_{B}$-normal form. Then, $n$ is in $U^{b}_{B}$-normal form, so by Lemma 7.7, ${\uparrow}n=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n$ is in $U^{c}_{C}$-normal form, which since $c={\rm Base}_{C}({\uparrow}n)$ implies that ${\uparrow}n$ is in $U_{C}$-normal form. Let $y\in{\rm Digit}_{c}({\uparrow}n)$. By Lemma 3.5(3), $y={\uparrow}x$ for some $x\in{\rm Digit}_{b}(n)$. Since $n$ is in $u_{B}$-normal form, $x$ is in $u_{B}$-normal form and $u_{B}(x)<u_{B}(n)$. By the induction hypothesis, $y={\uparrow}x$ is in $u_{C}$-normal form and
@@
where the first equality is by the induction hypothesis and we have already established the second. ∎
Next, our goal is to match up the fundamental sequences with values of the new ordinal interpretation. This is where greedy successors come in. To elucidate the relation between them and the fundamental sequences, let $B$ be a base hierarchy and $C=B^{+}$. Suppose that $B$ can ‘compute fundamental sequences’ for $n$ up to some $i\in\mathbb{N}$, in the sense that there is $m<n$ such that $u_{B}(m)=u_{B}(n)[i]$. The claim is that $C$ can then compute fundamental sequences for ${\uparrow}n$ up to $i+1$. The critical case is that where $\zeta=U_{B}(n)$ has uncountable cofinality. Then, $u_{B}(n)[i+1]=\psi(\zeta[u_{B}(n)[i]])=\psi(\zeta[u_{B}(m)])$. Intuitively, $\zeta[u_{B}(m)]$ is obtained by replacing the ‘rightmost’ occurrence of $\Omega$} by $u_{B}(m)$, e.g. if $\zeta={\Omega}^{\Omega}$, then $u_{B}(n)[i+1]=\psi({\Omega}^{u_{B}(m)})$. If $b={\rm Base}_{B}(n)$, this would be reflected by replacing the ‘rightmost’ occurrence of $b$ by $m$, so that if $n=b^{b}$, we would have $u_{B}(n)[i+1]=u_{B}(b^{m})$. But, we may not be able to do this, since we might have that $m>b$.
However, the situation changes on $C$. Since $\zeta$ has uncountable cofinality, it must be a multiple of $\Omega$}, so $n$ is a multiple of $b$ by Lemma 7.2(2); in other words, $n\in\partial B$. Using Proposition 5.2(2), we see that ${\uparrow}n=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n$, where $c>{\uparrow}(n-1)$, i.e. we added a base between ${\uparrow}(n-1)$ and ${\uparrow}n$, which by monotonicity, yields ${\uparrow}m<c$. But that means that ${\uparrow}m$ is now available as a $c$-digit, so we can indeed substitute it for the rightmost $c$ in ${\uparrow}n$, thus obtaining $c^{{\uparrow}m}<{\uparrow}n$ with $u_{C}(c^{{\uparrow}m})=u_{C}({\uparrow}n)[i+1]=u_{B}(n)[i+1]$. In the countable cofinality case, we simply replace the rightmost coefficient $a$ by $a^{\prime}$ with $u_{C}(a^{\prime})=u_{C}(a)[i+1]$, assuming inductively that $C$ can also compute the fundamental sequences for $a$. This idea of replacing the rightmost ocurrence of a base or coefficient to simulate the fundamental sequences is made precise in the next lemma.
@@
  1. (1)
    If $U(s)=\xi+1$, then $s>0$, $U(s-1)=\xi$, and s1s-1 is in $U^{b}_{B}$-normal form.
  2. (2)
    If $U(s)$ is a limit, suppose that $\iota<\tau(U(s))$ and $\Delta\in(0,\min B]$ are such that $\iota=u_{B}(c)$ for some $c\leq b-\Delta$ in $u_{B}$-normal form, and either $\iota$ is infinite, or for every $b$-digit $a$ of $s$, if $U(a)$ is a limit, then there is $a^{\prime}\leq a-\Delta$ in $u_{B}$-normal form such that $u_{B}(a^{\prime})=u_{B}(a)[\iota]$.
    Then, there is $s^{\prime}<s$ in $U^{b}_{B}$-normal form such that $U(s^{\prime})=U(s)[\iota]$ and $s^{\prime}\leq s-\Delta$.
Proof.
We prove both claims simultaneously by induction on $s+b$.
  1. Case 1
    ($s<b$). Then, $U(s)=u_{B}(s)<{\Omega}$ and thus has countable cofinality, so $\iota$ is finite. Consider the following sub-cases.
    1. Case 1.1
      ($U(s)$ is a successor). In this case, either $s<\min B$, $U(s)=s$, and U(s)[i]=s1U(s)[i]=s-1 by Lemma 6.5(1), so we can take $s^{\prime}=s-1$, or else $b^{\prime}:={\rm Base}_{B}(s)<b$ and we can again take $s^{\prime}=s-1$ by the induction hypothesis given that $s+b^{\prime}<s+b$.
    2. Case 1.2
      ($U(s)$ is a limit). By assumption, we can find $s^{\prime}\leq s-\Delta$ in $u_{B}$-normal form (which, since $s^{\prime}<b$, implies that $s^{\prime}$ is in $U^{b}_{B}$-normal form by definition) with $u_{B}(s^{\prime})=u_{B}(s)[\iota]$, and since $s^{\prime}<s<b$, we have that $U(s^{\prime})=u_{B}(s^{\prime})$, so